OpenStudy (anonymous):

Can someone help me with using the addition-subtraction method? x+y=7 x-y+3

OpenStudy (anonymous):

*x-y=3

OpenStudy (anonymous):

x+y=7 x-y=3 now add the two and you get 2x = 10 ( y and -y when added = 0) so x = 10/2 =5 now substitute x=5 in any of the given equation and get value of y

OpenStudy (anonymous):

Can you do that one by one with me in the equation form? I live better though example.

OpenStudy (anonymous):

Write the given equations one below the other. we see that in first one we have a positive y and in the second one we have negative y. when added they will cancel each other out. x+y=7 -- (I) x-y=3 -- (II) now add the two and you get ------ 2x = 10 x = 10/2 x = 5 Understood till this step ??

OpenStudy (anonymous):

OHHHHH! I get now I for some reason thought that one of the x's was negative and I was going crazy! Thanks so much!! :)

OpenStudy (anonymous):

u r welcome....☺ you know the steps after this...??

OpenStudy (anonymous):

Think you can help me with one more problem?

OpenStudy (anonymous):

@HappySoul